### v1

# MoCrop: Training Free Motion Guided Cropping for Efficient Video Action Recognition

Abstract
We introduce MoCrop, a motion-aware adaptive cropping module for efficient video action recognition in the compressed domain. MoCrop uses motion vectors that are available in H.264 video to locate motion-dense regions and produces a single clip-level crop that is applied to all I-frames at inference. The module is training free, adds no parameters, and can be plugged into diverse backbones. A lightweight pipeline that includes denoising & merge (DM), Monte Carlo sampling (MCS), and adaptive cropping (AC) via a motion-density submatrix search yields robust crops with negligible overhead. On UCF101, MoCrop improves accuracy or reduces compute. With ResNet-50, it delivers +3.5% Top-1 accuracy at equal FLOPs (attention setting), or +2.4% Top-1 accuracy with 26.5% fewer FLOPs (efficiency setting). Applied to CoViAR, it reaches 89.2% Top-1 accuracy at the original cost and 88.5% Top-1 accuracy while reducing compute from 11.6 to 8.5 GFLOPs. Consistent gains on MobileNet-V3, EfficientNet-B1, and Swin-B indicate strong generality and make MoCrop practical for real-time deployment in the compressed domain. Our code and models are available at https://github.com/microa/MoCrop.
Index Terms—  Motion-Aware Cropping, Video Action Recognition, Motion Vectors, Computational Efficiency, Plug-and-Play Module
1 Introduction
Mirroring the human ability to rapidly ignore distractors and attend to informative regions, modern vision systems use attention to highlight critical image areas [1].
Deep learning transformed video action recognition: from hand-engineered features to learned spatiotemporal representations [2, 3]. Two-Stream underscored motion via precomputed optical flow [4], while C3D and video transformers learn end to end features and reach strong accuracy [5, 6, 7]. The cost is substantial computation, motivating more efficient approaches.
A complementary route is to exploit the compressed domain [8]: codecs such as H.264 store I-frames and represent others using motion vectors (MV) and residuals [9, 10]. These cues emphasize dynamic regions that often coincide with actions [11]. Systems like CoViAR process I-frames, motion vectors, and residuals in separate streams [8], and other works refine motion vectors toward optical flow [12]. In effect, compression already performs a form of feature selection [13].
Motivated by this, we hypothesize that a simple, parameter-free cropping strategy can capture spatially salient regions (SSR) using motion vectors alone.
We introduce MoCrop, a lightweight, training-free module that selects key spatial regions from motion vectors. We apply Monte Carlo downsampling [14], accumulate sampled motion vectors on a spatial grid, and derive a single crop for the clip. MoCrop attaches at inference, requires no model changes or retraining, and contrasts with many attention mechanisms that are embedded and trained end to end [15, 7]. Empirically, it improves accuracy while lowering inference FLOPs.
Scope and assumptions. MoCrop assumes access to codec motion vectors and focuses on a single crop per clip for temporal stability. The design is orthogonal to backbone choice, training protocol, and test-time augmentation, and it can be enabled or disabled without retraining.
Contributions.
  • A training-free, plug-and-play MV-driven adaptive cropping module.
  • Accuracy and efficiency gains across common backbones.
  • A lightweight design using Monte Carlo downsampling on motion vectors to reduce computation while capturing discriminative regions.
2 Method
Fig. 1: MoCrop pipeline. Motion vectors identify spatially salient regions (SSR) to guide I-frame cropping; cropped frames are fed to the action recognition model.
MoCrop is a preprocessing module that uses motion vectors from compressed streams to locate spatially salient regions (SSR) and compute a single adaptive crop, focusing inference on informative regions. This reduces input resolution and FLOPs and often improves accuracy by suppressing background distraction.
2.1 Problem Formulation
Given decoded I-frames and raw motion vectors, let $V_{\rm raw}=\{(p_{k},\Delta p_{k})\}_{k=1}^{T}$ with $p_{k}=(x_{k},y_{k})$ the origin and $\Delta p_{k}$ the displacement. We seek a function $f$ that maps motion vectors to a clip-level box $B^{\ast}$:
$B^{\ast}=f(V_{\rm raw}),$
computed once per clip for temporal consistency. Each I-frame $I_{i}$ is then cropped by
$I_{i}^{\prime}=g(B^{\ast},I_{i}).$
2.2 Detailed Algorithm
MoCrop comprises three processing modules, DM (denoise & merge), MCS (Monte Carlo sampling), and AC (adaptive cropping). DM suppresses static/low-amplitude MVs and merges them across frames; MCS randomly subsamples the filtered MV to a fixed budget and counts them per cell to form the motion-density map (MD map); AC aggregates MD on a grid, searches a high-score rectangle near a target area ratio, and crops the clip accordingly.
2.2.1 Stage 1: Denoise-and-Merge (DM)
Extraction and Filtering. Let $V_{\rm raw}=\{(p_{k},\Delta p_{k})\}_{k=1}^{T}$. We suppress static/near-static vectors:
$V_{\rm filtered}=\{(p,\Delta p)\in V_{\rm raw}:\ \|\Delta p\|_{2}>\epsilon\},$
with small threshold $\epsilon$.
2.2.2 Stage 2: Monte Carlo Sampling & MD Map
Monte Carlo Sampling (MCS). For efficiency when $|V_{\rm filtered}|$ is large, we uniformly sample a subset $V_{\rm sampled}\subseteq V_{\rm filtered}$ of size $|V_{\rm sampled}|=\min(N,|V_{\rm filtered}|)$, preserving spatial distribution while reducing points.
MD Map. We discretize a frame of size $(W_{\rm video},H_{\rm video})$ into an $h\times w$ grid. Cell $C_{ij}$ spans
$C_{ij}=\Big[j\tfrac{W_{\rm video}}{w},(j+1)\tfrac{W_{\rm video}}{w}\Big)\times\Big[i\tfrac{H_{\rm video}}{h},(i+1)\tfrac{H_{\rm video}}{h}\Big).$
The map $M$ counts sampled motion vectors per cell:
$M_{ij}=\sum_{(p_{k},\cdot)\in V_{\rm sampled}}I(p_{k}\in C_{ij}),$
where $I(\cdot)$ is the indicator; larger $M_{ij}$ indicates stronger motion.
2.2.3 Stage 3: Submatrix Search and Box Generation (AC)
We find the submatrix with maximal motion mass:
$(i^{\ast},j^{\ast},H_{s}^{\ast},W_{s}^{\ast})=\mathop{\rm arg\,max}_{i,j,H_{s},W_{s}}\ \sum_{p=i}^{i+H_{s}-1}\ \sum_{q=j}^{j+W_{s}-1}M_{pq}$
subject to
$\begin{array}[]{l}1\leq H_{s}\leq h,\quad 1\leq W_{s}\leq w,\\[2.0pt] 0\leq i\leq h-H_{s},\quad 0\leq j\leq w-W_{s},\\[2.0pt] (1-\delta)\alpha hw\leq H_{s}W_{s}\leq(1+\delta)\alpha hw,\end{array}$
with target area ratio $\alpha\in(0,1]$ and tolerance $\delta$ (e.g., $0.1$). We convert grid parameters to a normalized YOLO [16] box $(c_{x},c_{y},w_{b},h_{b})$:
$c_{x}=\frac{j^{\ast}+W_{s}^{\ast}/2}{w},\ c_{y}=\frac{i^{\ast}+H_{s}^{\ast}/2}{h},\ w_{b}=\frac{W_{s}^{\ast}}{w},\ h_{b}=\frac{H_{s}^{\ast}}{h}.$
Here $i,j,H_{s},W_{s}\in\mathbb{N}$ denote integer grid indices and sizes.
2.2.4 Stage 4: Cropping I-frames
We map the normalized box to pixel coordinates $(x_{1},y_{1},x_{2},y_{2})$:
$\begin{array}[]{l}x_{1}=\max\Big(0,\ \lfloor(c_{x}-w_{b}/2)\,W_{\rm video}\rfloor\Big),\\[3.0pt] y_{1}=\max\Big(0,\ \lfloor(c_{y}-h_{b}/2)\,H_{\rm video}\rfloor\Big),\\[3.0pt] x_{2}=\min\Big(W_{\rm video},\ \lfloor(c_{x}+w_{b}/2)\,W_{\rm video}\rfloor\Big),\\[3.0pt] y_{2}=\min\Big(H_{\rm video},\ \lfloor(c_{y}+h_{b}/2)\,H_{\rm video}\rfloor\Big).\end{array}$
and extract
$I_{\rm crop}=I[y_{1}:y_{2},\ x_{1}:x_{2}].$
The same box is applied to all I-frames in the clip, yielding an efficient, training-free, plug-and-play preprocessor.
Components. DM: denoise & temporal merge of motion vectors; MCS: Monte Carlo sampling on filtered MVs to estimate a motion-density (MD) map on an $h{\times}w$ grid; AC: grid-wise aggregation and rectangle search over MD to produce a single clip-level crop (Fig. 1).
2.3 Complexity and Practicality
Complexity. DM scans raw motion vectors once to threshold and (optionally) merge across frames, i.e., $O(|V_{\rm raw}|)$ time.
Building the motion-density map costs $O(|V_{\rm sampled}|+hw)$ for counting on an $h\times w$ grid. For submatrix search, our prototype directly sums each candidate window on an $R{\times}C$ grid (R=hR{=}h, C=wC{=}w): the worst-case time is $O(R^{3}C^{3})$ (looping over $(H_{s},W_{s})$ and positions with $O(H_{s}W_{s})$ sum per window). In practice $R{\times}C\!\leq\!48$ (e.g., $6{\times}8$), so this overhead is negligible compared with backbone inference. Optionally, a 2D integral image reduces each window score to $O(1)$, yielding $O(\#\text{cands})$ overall; when $(H_{s},W_{s})$ are fixed, a sliding-window implementation is $O(hw)$ total.
Hyperparameters. $\epsilon$ can be set by a small absolute magnitude or by a low percentile of $\|\Delta p\|_{2}$; $N$ trades variance and cost and is robust across a range; $(h,w)$ can be chosen proportional to frame size and codec macroblocks; $(\alpha,\delta)$ control the crop area and tolerance. For strong camera motion, subtracting a global estimate (for example, the median motion vector) stabilizes the map. When the map is flat, falling back to a center crop is a safe default.
3 Experiments
3.1 Dataset and Metrics
All experiments are conducted on UCF101, a widely used benchmark for human action recognition featuring 13,320 videos across 101 action classes [17]. Given that our approach is designed for efficiency, we evaluate performance by jointly considering Top-1 accuracy and computational cost in GFLOPs.
Table 1: Demonstration that MoCrop provides dual advantages as a plug-and-play module on modern backbones using the UCF101 dataset. The table highlights the performance gains in both the accuracy-focused Attention Setting and the cost-saving Efficiency Setting.
Backbone Dataset Variant Top-1 Acc. (%) GFLOPs $\Delta$ Acc. (%) $\Delta$ FLOPs (%)
ResNet-50 [18] UCF101 (A) Baseline (224px) 80.1 4.11 - -
(B) + MoCrop (Efficiency, 192px) 82.5 3.02 +2.4 -26.5
(C) + MoCrop (Attention, 224px) 83.6 4.11 +3.5 0.0
MobileNet-V3 [19] UCF101 (A) Baseline (224px) 78.3 0.22 - -
(B) + MoCrop (Efficiency, 192px) 79.3 0.17 +1.0 -22.7
(C) + MoCrop (Attention, 224px) 80.2 0.22 +1.9 0.0
EfficientNet-B1 [20] UCF101 (A) Baseline (224px) 82.1 0.59 - -
(B) + MoCrop (Efficiency, 192px) 83.6 0.43 +1.5 -27.1
(C) + MoCrop (Attention, 224px) 84.5 0.59 +2.4 0.0
Swin-B [6] UCF101 (A) Baseline (224px) 87.3 15.5 - -
(B) + MoCrop (Efficiency, 192px) 87.2 12.6 -0.1 -18.7
(C) + MoCrop (Attention, 224px) 87.5 15.5 +0.2 0.0
3.2 Experimental Setup
We evaluate MoCrop on the official UCF101 split. Following standard practice [8], we sample three frames per video for training and use typical data augmentation. At inference, we freeze the model weights and evaluate two main settings: Baseline: The model trained on UCF101 is evaluated on the standard test set with a 224px input. MoCrop: The same model is evaluated after each video is pre-processed by MoCrop. We test two variants:
  • Efficiency Setting: MoCrop crops the video, which is then resized to a smaller resolution (192px) to reduce FLOPs while aiming to maintain or improve accuracy.
  • Attention Setting: MoCrop crops the video, which is resized back to the original resolution (224px). This isolates the accuracy improvement gained by focusing on salient regions, without changing the computational cost.
3.3 Performance and Efficiency Analysis
Main Results.
Table 1 demonstrates the powerful dual advantages of MoCrop as a plug-and-play module for various modern backbones. In the Attention Setting, where FLOPs are held constant, MoCrop consistently boosts performance by forcing the model to concentrate on motion-rich areas. The most significant gain is seen with ResNet-50 [18], which improves by a remarkable +3.5% Top-1 accuracy. Other models like EfficientNet-B1 [20] (+2.4%) and MobileNet-V3 [19] (+1.9%) also show substantial improvements.
In the Efficiency Setting, MoCrop proves its ability to reduce computational costs and often enhances accuracy. For instance, ResNet-50 sees a +2.4% accuracy increase while FLOPs are cut by a massive 26.5%. Even for the heavy Swin-B Transformer [6], where accuracy gains are harder to achieve, our method reduces the computational load by 18.7% with only a negligible 0.1% accuracy dip. These results confirm that MoCrop offers an exceptional and highly flexible trade-off between accuracy and efficiency.
3.4 Ablation Studies
To validate the contributions of the internal components of MoCrop and its effectiveness as a general module, we present a series of ablation studies in Table 2. The studies are organized into three parts: (a) an incremental analysis of our module’s components, (b) a comparison against standard cropping methods, and (c) an application of MoCrop to enhance a prior state-of-the-art method.
Incremental Component Analysis. Using ResNet-18 [18] as a base, we find that each component of MoCrop contributes to the final performance. Simply adding Adaptive Cropping (AC) improves accuracy from 79.3% to 80.6%. The addition of Denoise & Merge (DM) and Monte Carlo Sampling (MCS; MD map) further pushes the accuracy to a final 81.4%. This confirms that all parts of our design are integral to its success.
Comparison with Cropping Strategies. MoCrop significantly outperforms standard cropping techniques. While 90% Random and Center crops offer a slight improvement over the baseline, our adaptive method achieves a much higher accuracy of 81.4% with fewer FLOPs. This demonstrates the superiority of using motion information to guide cropping decisions over static, fixed-ratio approaches.
Enhancing Prior Works. To show its versatility, we applied MoCrop to CoViAR, a well-known compressed-domain method. The results are compelling. In the efficiency setting, MoCrop improves CoViAR’s accuracy to 88.5% while reducing its FLOPs from 11.6 GFLOPs to just 8.5 GFLOPs. In the attention setting, it boosts the accuracy to an impressive 89.2%. This shows that MoCrop is not just a standalone module but can also serve as a powerful enhancer for other methods in the field.
Table 2: Ablation studies and comparative analysis on ResNet-18 and CoViAR. MoCrop shows clear benefits over baselines and standard techniques, and effectively enhances prior state-of-the-art methods.
Setup Method Acc. (%) GFLOPs
(a) Ablation on MoCrop Components (on ResNet-18 [18])
ResNet-18 [18] Baseline 79.3 1.82
+ AC 80.6 1.54
+ AC + DM 81.3 1.54
+ AC + DM + MCS (Full) 81.4 1.54
(b) Comparison with Cropping Strategies (on ResNet-18)
ResNet-18 Baseline 79.3 1.82
Random Crop (90%) 80.4 1.73
Center Crop (90%) 80.2 1.73
MoCrop (Adaptive) 81.4 1.54
(c) Enhancing Prior Works (on UCF101, I-frames only)
CoViAR [8] Baseline (Reported in [8]) 87.7 11.6
+ MoCrop (Efficiency, 192px) 88.5 8.5
+ MoCrop (Attention, 224px) 89.2 11.6
(a) PullUps
(b) ApplyEyeMakeup
(c) JumpRope
(d) Archery
(e) BalanceBeam
(f) HighJump
Fig. 2: Visualization of MoCrop (green box) vs. center cropping (red box), alongside motion-density (MD) maps. Each subfigure presents two views in order: RGB frame with boxes (left) and the MD map (right), shown across six UCF101 classes.
3.5 Visualization
Fig. 2 presents qualitative examples of adaptive cropping with MoCrop. For each class, the left panel shows an RGB frame with the green box predicted by MoCrop and the red fixed center-crop box; the right panel shows the corresponding motion-density (MD) map. Across six UCF101 classes (PullUps, ApplyEyeMakeup, JumpRope, Archery, BalanceBeam, and HighJump), MoCrop places a tighter crop around the actor and the regions with salient motion (e.g., the torso during pull-ups or the take-off area in high jump). In contrast, the center crop often includes large static background or truncates the subject (notably in Archery and BalanceBeam). The MD maps on the right concentrate energy near the actor and align with the green crop, illustrating the signal that guides our policy.
4 Discussion and Conclusion
We present MoCrop, a lightweight, training-free module using compressed-domain motion vectors to generate a single, clip-level crop of motion-dense regions. By focusing inference, MoCrop reduces resolution and FLOPs, often boosting accuracy, and integrates with existing backbones without retraining.
When it helps. MoCrop is most effective when actions coincide with localized motion and when background clutter dilutes fixed crops. The map-based formulation is robust to moderate noise after filtering and sampling, and it generalizes across backbones since it operates as a preprocessing step.
Limitations. MoCrop relies on codec-provided motion vectors (e.g., H.264) and is less informative for raw videos unless motion estimation is run in parallel. This dependency is acceptable since most datasets and streaming pipelines use compressed video. Large camera motion or scene-wide dynamics can bias the density map; mitigations include global-motion suppression, temporal smoothing, and combining motion vector magnitude with residual cues.
Future work. We plan to extend MoCrop to diagonally elongated or multi-region actions via multi-box proposals, explore codec-agnostic motion signals, and validate on larger datasets and backbones. We will also study learned selection over MoCrop proposals while keeping inference cost low.
In summary, MoCrop offers a practical plug-and-play path to efficiency and accuracy in compressed-domain video understanding, requiring no additional parameters and minimal compute overhead.